# bibdata.bib

### v1

@@ IV-B1 Theoretical Foundation
Logic simulation can obtain a lower bound of the maximum error caused by a candidate LAC $l_{\textit{cand}}$. Consider the circuit after applying LAC $l_{\textit{cand}}$, denoted as $\hat{G}_{\textit{cand}}$, and denote the set of all PI patterns as $\mathbb{B}^{I}$. Under a set of $M$ sampled PI patterns, denoted as $\mathbb{S}\subseteq\mathbb{B}^{I}$, we simulate the accurate circuit $G$ and the approximate circuit $\hat{G}_{\textit{cand}}$. We obtain the maximum deviation between the POs of circuit $G$ and $\hat{G}_{\textit{cand}}$ over input patterns in $\mathbb{S}$ and denote it as ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$. Then, clearly, ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$ is a lower bound on the maximum error caused by $l_{\textit{cand}}$ , i.e., ${\textit{MaxError}}(\hat{G}_{\textit{cand}})$, since the calculation of ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$ does not consider all PI patterns. In summary, we have
${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})\leq{\textit{MaxError}}(\hat{G}_{\textit{cand}}).$ (4)
Eq. (4) is the basis of the simulation-guided LAC pruning. If simulation finds that ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})>B$, then we have ${\textit{MaxError}}(\hat{G}_{\textit{cand}})>B$, indicating that the LAC $l_{\textit{cand}}$ is invalid and should be discarded. Otherwise, if ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})\leq B$, then the LAC $l_{\textit{cand}}$ is retained in the set of remaining LACs $L_{\textit{rem}}$. In this case, the actual maximum error caused by $l_{\textit{cand}}$ may still exceed $B$ under some unsampled PI patterns, and SAT-based formal error checking will be performed to check whether $l_{\textit{cand}}$ really satisfies the maximum error constraint or not. In our implementation, we first choose a small number of $M_{\textit{small}}$ random PI patterns to roughly prune the invalid LACs, followed by a larger number of $M$ PI patterns to obtain a tighter lower bound of the maximum error to further prune the invalid LACs. The LACs not pruned by the two-round simulation are retained in $L_{\textit{rem}}$.
@@ IV-B2 Efficient Computation of the Maximum Error Lower Bounds
The theoretical foundation of the simulation-guided LAC pruning in Eq. (4) requires computing ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$ for each candidate LAC $l_{\textit{cand}}$. To compute ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$ for each candidate LAC $l_{\textit{cand}}$, a naive way is to apply $l_{\textit{cand}}$ to the current approximate circuit $\hat{G}$ and obtain a new approximate circuit $\hat{G}_{\textit{cand}}$. After simulating the accurate circuit $G$ and the approximate circuit $\hat{G}_{\textit{cand}}$ under the PI patterns in $\mathbb{S}$, ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$ can be obtained. This method is straightforward but slow, requiring $O(|L_{\textit{cand}}|)$ simulation runs, where $|L_{\textit{cand}}|$ is the number of candidate LACs.
Instead of using the naive method, we accelerate the computation of all ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$’s based on the change propagation matrix (CPM) proposed in [su2022vecbee]. The CPM $P$ for the current approximate circuit $\hat{G}$ is a three-dimensional 0-1 matrix of size $M\times N\times O$, where $M$ is the number of simulation patterns, $N$ is the number of functional nodes in the circuit, and $O$ is the number of POs in the circuit. Each entry in the CPM is indexed as $P[i,n,\hat{y}_{k}]$, where $1\!\leq\!i\!\leq\!M$ represents the $i$-th simulation pattern, $n$ is a functional node in the circuit, and $\hat{y}_{k}$ ($1\!\leq\!k\!\leq\!O$) is the $k$-th PO in $\hat{G}$. The entry $P[i,n,\hat{y}_{k}]$ evaluates the impact of the change in $n$’s value on $\hat{y}_{k}$. Specifically, $P[i,n,\hat{y}_{k}]=1$ indicates that a flip of $n$’s value will cause a flip of $y_{k}$’s value under the $i$-th pattern, while $P[i,n,\hat{y}_{k}]=0$ means that $\hat{y}_{k}$’s value keeps unchanged after a flip of $n$’s value under the $i$-th pattern. To compute $P[i,n,\hat{y}_{k}]$, under the $i$-th pattern we can flip $n$’s value, update the values of all $n$’s TFOs and POs using $n$’s new value, and then check whether $\hat{y}_{k}$’s value changes. If it changes, then $P[i,n,\hat{y}_{k}]$ is 1; otherwise, $P[i,n,\hat{y}_{k}]$ is 0. We apply the above process to each funtional node $n$ in the circuit to obtain its CPM entries. Thus, computing the CPM for all functional nodes requires $O(N)$ simulation runs, where $N$ is the number of functional nodes.
@@
CPM can be used to efficiently compute the PO values after applying each candidate LAC, and hence the lower bound on the maximum error of the candidate LAC. For the single-output LACs considered in this work, each of them, denoted as $l_{\textit{cand}}$, can be modeled as replacing an existing node $n$ in the circuit with a new node $n^{\prime}$. After applying $l_{\textit{cand}}$ to the current approximate circuit $\hat{G}$, we can obtain the new value of the $k$-th PO under the $i$-th pattern, denoted as $y^{\prime}_{k}[i]$, as follows:
$\begin{split}y^{\prime}_{k}[i]&=\hat{y}_{k}[i]\oplus{\textit{Impact}}[i,l_{\textit{cand}},\hat{y}_{k}]\\ &=\hat{y}_{k}[i]\oplus\left((n[i]\oplus n^{\prime}[i])\land P[i,n,\hat{y}_{k}]\right),\\ \end{split}$ (5)
@@
$D\left(\bm{y}[i],\bm{\hat{y}}[i]\right)=\left|(2y_{1}[i]+y_{0}[i])-(2y^{\prime}_{1}[i]+y^{\prime}_{0}[i])\right|=1.\$
@@
Using the CPM-based method to compute the lower bounds on the maximum errors of all candidate LACs, the main computational effort lies in constructing the CPM. As mentioned above, obtaining the CPM for the current approximate circuit $\hat{G}$ requires only $O(N)$ simulation runs, where $N$ is the number of functional nodes in $\hat{G}$. Compared with the naive method with $O(|L_{\textit{cand}}|)$ simulation runs, the CPM-based method is much more efficient, since $N$ is usually much smaller than $|L_{\textit{cand}}|$. After obtaining the lower bounds on the maximum errors of all candidate LACs, the LAC pruning step can efficiently filter out the invalid LACs that violate the maximum error constraint based on Eq. (4) and return the set of remaining LACs $L_{\textit{rem}}$ to be further checked by the promising LAC selection and application step.
@@ IV-C Promising LAC Selection and Application Based on Simulation-Guided SAT Solving
As shown in Fig. 2, the promising LAC selection and application step is responsible for checking the validity of the remaining LACs in $L_{\textit{rem}}$ after the simulation-based pruning, selecting a subset of promising LACs $L_{\textit{prom}}$ to reduce the circuit area as much as possible, and applying all LACs in $L_{\textit{prom}}$ to simplify the current approximate circuit $\hat{G}$. In the following parts, we will first formulate the LAC selection problem in Section IV-C1 and then introduce a greedy LAC selection strategy supported by simulation-guided SAT solving in Section IV-C2. Finally, we will discuss the order of checking LACs in Section IV-C3, which significantly affects the quality of the final approximate circuit and the efficiency of the promising LAC selection and application step.
@@ IV-C1 Formulation of the LAC Selection Problem
This problem is similar to the one in the MUSCAT method [witschen2022muscat], which is a state-of-the-art ALS method under the maximum error constraint. However, the formulation of MUSCAT only considers the constant LACs that replace a signal by a constant 0 or 1, while ours considers arbitrary single-output LACs, To solve this NP-hard combinatorial optimization problem, MUSCAT converts the problem into a MUS problem and solves it using MUS solvers, which is time-consuming and limits its scalability. Although MUSCAT sets a time limit for the MUS solving to enhance scalability, the time limit leads to suboptimal approximate circuits. In our work, we propose a greedy selection strategy to obtain a good solution efficiently, which will be introduced next.
@@ IV-C2 Greedy LAC Selection Strategy Supported by Simulation-Guided SAT Solving
  • UNSAT: This indicates that the maximum error caused by LAC $l_{j}$ is no larger than $B$. In this case, LAC $l_{j}$ is valid, and we apply it to the circuit $G_{j-1}$, obtaining the resulting circuit $G_{j}$. Examples of this case are $l_{1}$ and $l_{3}$ in Fig. 4. Note that in order to satisfy the LAC conflict constraint, after applying $l_{j}$, we remove all LACs that affect the same node as $l_{j}$ from $L_{\textit{rem}}$.
@@
Figure 4: Greedy-based promising LAC selection strategy supported by simulation-guided SAT solving.
@@
We propose a reuse mechanism of the counter-example patterns, as shown in the right part of Fig. 4. When checking the maximum error of LAC $l_{j}$, before solving the SAT problem for $l_{j}$, we perform logic simulation using the counter-example patterns generated by $l_{s}$’s ($s<j$) to quickly check the maximum error caused by $l_{j}$. If simulation finds that $l_{j}$ is invalid, then we directly skip the LAC without solving the SAT problem. Otherwise, we still need to solve the SAT problem for $l_{j}$ to check its validity. By guiding the SAT solving with the simulation using the counter-example patterns, we can skip unnecessary SAT solving and improve the efficiency of the LAC selection and application step. Furthermore, we also update the simulation patterns by storing the counter-example patterns. In this way, the stored counter-example patterns will guide the future iterations of the ALS flow in their simulation-based LAC pruning and promising LAC selection and application steps. In practice, this method reduces the number of SAT problems to be solved in the ALS flow and accelerates the ALS flow.
@@ IV-C3 The Order of Checking and Selecting LACs
We propose to sort the LACs by the ascending order of their maximum error lower bounds, obtained from logic simulation results. The motivation is that only the top $K$ LACs are kept after the sorting, and we want to keep the valid LACs in the top $K$ LACs as many as possible so that the circuit area can be effectively reduced by applying these valid LACs. Given that a LAC with a small maximum error lower bound is more likely to be valid, sorting the LACs by the lower bounds can increase the probability of keeping valid LACs in the top $K$ LACs.
Besides considering the maximum error lower bound of each LAC as the primary sorting criterion, our sorting strategy also considers the reduced area of each LAC. Specifically, for LACs with the same maximum error lower bound, we sort them by the reduced area in descending order, since LACs reducing more area can simplify the circuit more effectively and should be considered first.
@@ V-A Experimental Setup
We implement the proposed ALS flow under the maximum error constraint in C++ and test it on a single core of an AMD Ryzen 9 5900X processor with 64GB RAM. Our flow is developed upon a state-of-the-art open-source logic synthesis and verification system, ABC [berkeley2024abc]. The flow also integrates CryptoMiniSat [soos2009extending], a modern SAT solver with rich features and a simple interface. To avoid long runtime of SAT solving, we set a maximum conflict number of $2^{18}$ for each SAT problem.
@@
In all experiments, the original circuits are first converted into AIGs and then simplified by our flow. The reason for using AIGs is that many AIG-based logic synthesis works [riener2019scalable, calvino2022versatile] and ALS works [meng2023hedals, lee2023approximate] have shown significant advantages in reducing hardware cost, particularly for CMOS technologies. The standard cell library used in our experiments is the Nangate 45nm library [nangate2022nangate]. For the traditional logic synthesis in the last step of our flow shown in Fig. 2, we use a script that performs “resyn2rs” for 3 times to optimize the AIG, followed by “dch; amap” for area-oriented technology mapping. Moreover, two types of LACs, the constant [shin2011new] and the SASIMI LACs [venkataramani2013substitute], are used to approximately simplify the circuits in our implementation, although our flow can be extended to support other single-output LACs. For the simulation-based LAC pruning (see Section IV-B1), we first use $M_{\textit{small}}=2^{10}$ simulation patterns to quickly filter out large-error circuits, and then use $M=2^{13}$ simulation patterns for a more fine-grained filtering. For the parameter $K$ in the sorting strategy in Section IV-C3, we set $K=100$ in all experiments. To build the error miter in Fig. 1 for maximum error checking, we use Verilog to describe the error miter and then use Yosys [yosys2025yosys] and ABC to synthesize the Verilog description.
@@
To evaluate the hardware cost of a circuit, we use area ratio (the area of the approximate circuit over that of the accurate one) and delay ratio (the delay of the approximate circuit over that of the accurate one). Smaller ratios are preferred due to more reduction in area and delay. To evaluate the accuracy of circuits, two different maximum error metrics, MaxED and MaxHD, are considered in our experiments. Note that the focus of this work is ALS under the maximum error constraint, so we do not compare it with other ALS methods under the average error constraints, such as [lee2023approximate] and [ma2021approximate]. For all generated the approximate circuits, the error miter in Fig. 1 is used to formally verify that the maximum errors of the circuits satisfy the given error bounds.
@@
TABLE I: Experimental benchmarks. Area and delay are measured by mapping the AIGs into the Nangate 45nm library.
@@
Size Depth Area/$\mu m^{2}$ Delay/ns
Used in MECALS [meng2023mecals] (arithmetic) absdiff 16/8 141 14 87.3 0.42
add8 16/9 66 10 42.0 0.36
add32 64/33 252 64 184.6 1.84
binsqrd 16/18 1562 50 1052.3 1.53
buttfly 32/34 265 48 170.5 1.01
mac 12/8 145 20 92.8 0.60
mult8 16/16 649 40 435.4 1.26
mult16 32/32 1981 72 1418.8 1.98
mult32 64/64 8340 53 5723.3 1.87
EPFL arithmetic* add128 256/129 1297 28 933.4 0.96
bar 135/128 2688 14 1267.8 0.92
log2 32/32 38540 419 21480.6 14.26
max 512/130 2686 549 1646.3 15.81
mult64 128/128 33242 326 16447.3 9.47
sine 24/25 7044 180 4112.1 5.93
sqrt 128/64 21951 4591 13464.1 216.92
square 64/128 20030 296 12801.8 7.96
ISCAS85 c880 60/26 313 22 198.2 0.59
c1355 41/32 390 16 235.9 0.56
c1908 33/25 367 25 229.6 0.86
c2670 233/140 579 17 385.2 0.68
c3540 50/22 937 32 521.1 1.02
c5315 178/123 1306 28 720.3 0.72
c7552 207/108 1469 26 903.6 1.43

### v2

@@ IV-B1 Theoretical Foundation
Given a candidate LAC $l_{\textit{cand}}$, logic simulation over a subset of PI patterns yields the maximum observable error under these patterns, denoted as ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$. Since the simulation does not cover all PI patterns, this value is a lower bound on the real maximum error caused by $l_{\textit{cand}}$, i.e.,
${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})\leq{\textit{MaxError}}_{\textit{real}}(l_{\textit{cand}}).$ (4)
This property forms the basis for the simulation-guided LAC pruning. If simulation finds that ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})\!>\!B$, then $l_{\textit{cand}}$ is guaranteed to violate the maximum error constraint and is discarded. In our implementation, we first simulate with a small number of $M_{\textit{small}}$ random PI patterns to roughly prune the invalid LACs, followed by a larger number of $M$ PI patterns to obtain a tighter lower bound of the maximum error for further pruning. The LACs not pruned by the two-round simulation are retained in $L_{\textit{rem}}$.
@@ IV-B2 Efficient Computation of the Maximum Error Lower Bounds
The theoretical foundation of the simulation-guided LAC pruning in Eq. (4) requires computing ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$ for each candidate LAC $l_{\textit{cand}}$. To compute ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$ for each candidate LAC $l_{\textit{cand}}$, a naive way is to apply $l_{\textit{cand}}$ to the current approximate circuit $\hat{G}$ and obtain a new approximate circuit $\hat{G}_{\textit{cand}}$. After simulating the accurate circuit $G$ and the approximate circuit $\hat{G}_{\textit{cand}}$ under some sampled PI patterns, ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$ can be obtained. This method is straightforward but slow, requiring $O(|L_{\textit{cand}}|)$ simulation runs, where $|L_{\textit{cand}}|$ is the number of candidate LACs.
Instead of using the naive method, we accelerate the computation of all ${\textit{MaxError}}_{\textit{LB}}(l_{\textit{cand}})$’s based on the change propagation matrix (CPM) proposed in [33]. The CPM $P$ for the current approximate circuit $\hat{G}$ is a three-dimensional 0-1 matrix of size $M\times N\times O$, where $M$ is the number of simulation patterns, $N$ is the number of functional nodes in the circuit, and $O$ is the number of POs in the circuit. Each entry in the CPM is indexed as $P[i,n,\hat{y}_{k}]$, where $1\!\leq\!i\!\leq\!M$ represents the $i$-th simulation pattern, $n$ is a functional node in the circuit, and $\hat{y}_{k}$ ($1\!\leq\!k\!\leq\!O$) is the $k$-th PO in $\hat{G}$. The entry $P[i,n,\hat{y}_{k}]$ evaluates the impact of the change in $n$’s value on $\hat{y}_{k}$. Specifically, $P[i,n,\hat{y}_{k}]=1$ indicates that a flip of $n$’s value will cause a flip of $y_{k}$’s value under the $i$-th pattern, while $P[i,n,\hat{y}_{k}]=0$ means that $\hat{y}_{k}$’s value keeps unchanged after a flip of $n$’s value under the $i$-th pattern. To compute $P[i,n,\hat{y}_{k}]$, under the $i$-th pattern we can flip $n$’s value, update the values of all $n$’s TFOs and POs using $n$’s new value, and then check whether $\hat{y}_{k}$’s value changes. If it changes, then $P[i,n,\hat{y}_{k}]$ is 1; otherwise, $P[i,n,\hat{y}_{k}]$ is 0. We apply the above process to each functional node $n$ in the circuit to obtain its CPM entries. Thus, computing the CPM for all functional nodes requires $O(N)$ simulation runs, where $N$ is the number of functional nodes.
@@
CPM can be used to efficiently compute the PO values after applying each candidate LAC, and hence the lower bound on the maximum error of the candidate LAC. For the single-output LACs considered in this work, each of them, denoted as $l_{\textit{cand}}$, can be modeled as replacing an existing node $n$ in the circuit with a new node $n^{\prime}$. After applying $l_{\textit{cand}}$ to the current approximate circuit $\hat{G}$, we can obtain the new value of the $k$-th PO under the $i$-th pattern, denoted as $y^{\prime}_{k}[i]$, as follows: [^1] Eq. (5) is an intermediate result derived from [33]. However, [33] studies the ALS problem under the average error constraint. Here, we extend its use to the ALS problem under the maximum error constraint, where it serves as a key building block of our simulation-guided ALS flow.
[…]
@@
$D\left(\bm{y}[i],\bm{\hat{y}}[i]\right)=\left|(2y_{1}[i]+y_{0}[i])-(2y^{\prime}_{1}[i]+y^{\prime}_{0}[i])\right|=1.\\$
@@
Using the CPM-based method to compute the lower bounds on the maximum errors of all candidate LACs, the main computational effort lies in constructing the CPM. As mentioned above, obtaining the CPM for the current approximate circuit $\hat{G}$ requires only $O(N)$ simulation runs, where $N$ is the number of functional nodes in $\hat{G}$. Compared with the naive method with $O(|L_{\textit{cand}}|)$ simulation runs, the CPM-based method is much more efficient, as $N$ is usually much smaller than $|L_{\textit{cand}}|$. After obtaining the lower bounds on the maximum errors of all candidate LACs, the LAC pruning step can efficiently filter out the invalid LACs that violate the maximum error constraint based on Eq. (4) and return the set of remaining LACs $L_{\textit{rem}}$ to be further checked by the promising LAC selection and application step.
@@ IV-C Promising LAC Selection and Application Based on Simulation-Guided SAT Solving
As shown in Fig. 2, the promising LAC selection and application step is responsible for checking the validity of the remaining LACs in $L_{\textit{rem}}$ after the simulation-based pruning, selecting a subset of promising LACs $L_{\textit{prom}}$ to reduce the circuit area as much as possible, and applying all LACs in $L_{\textit{prom}}$ to simplify the current approximate circuit $\hat{G}$.
In the following parts, we will first formulate the LAC selection problem in Section IV-C1 and then introduce a greedy LAC selection strategy supported by simulation-guided SAT solving in Section IV-C2. Finally, we will discuss the order of checking LACs in Section IV-C3, which significantly affects the quality of the final approximate circuit and the efficiency of the promising LAC selection and application step.
@@ IV-C1 Formulation of the LAC Selection Problem
Note that our framework supports circuits represented in any graph format, such as AIGs and gate netlists. For AIGs, circuit area is estimated by counting the number of AND nodes, while for gate netlists, it is obtained by summing the areas of all gates.
Our LAC selection problem is similar to the one in the MUSCAT method [42], which is a state-of-the-art ALS method under the maximum error constraint. However, the formulation of MUSCAT only considers the constant LACs that replace a signal by a constant 0 or 1, while ours considers arbitrary single-output LACs. To solve this NP-hard combinatorial optimization problem, MUSCAT converts the problem into a MUS problem and solves it using MUS solvers, which is time-consuming and limits its scalability. Although MUSCAT sets a time limit for the MUS solving to enhance the scalability, the time limit leads to suboptimal approximate circuits. In our work, we propose a greedy selection strategy to obtain a good solution efficiently, which will be introduced next.
@@ IV-C2 Greedy LAC Selection Strategy Supported by Simulation-Guided SAT Solving
  • UNSAT: This indicates that the maximum error caused by LAC $l_{j}$ is no larger than $B$. In this case, LAC $l_{j}$ is valid, and we apply it to the circuit $G_{j-1}$, obtaining the resulting circuit $G_{j}$. Examples of this case are $l_{1}$ and $l_{3}$ in Fig. 4. Note that in order to satisfy the LAC conflict constraint, after applying $l_{j}$, we remove all LACs that affect the same node as $l_{j}$ from the top $K$ LACs.
@@
Figure 4: Greedy promising LAC selection strategy supported by simulation-guided SAT solving.
@@
We propose a reuse mechanism of the counter-example patterns, as shown in the right part of Fig. 4. When examining LAC $l_{j}$, before solving the SAT problem for $l_{j}$, we perform logic simulation using the counter-example patterns generated by $l_{s}$’s ($s<j$) to quickly check whether $l_{j}$ already violates the maximum error bound under these patterns. If simulation finds that $l_{j}$ is invalid, then we directly skip the LAC without solving the SAT problem. Otherwise, we still need to solve the SAT problem for $l_{j}$ to check its validity. By guiding the SAT solving with the simulation using the counter-example patterns, we can skip unnecessary SAT solving and improve the efficiency of the LAC selection and application step. Furthermore, we also update the simulation patterns by storing the counter-example patterns. In this way, the stored counter-example patterns will guide the future iterations of the ALS flow in their simulation-based LAC pruning and promising LAC selection and application steps. In practice, this method reduces the number of SAT problems to be solved in the ALS flow and accelerates the ALS flow.
@@ IV-C3 The Order of Checking and Selecting LACs
We propose a sorting strategy based on two criteria: 1) the maximum error lower bound (primary) and 2) the estimated area reduction of the LAC (secondary). First, the LACs are sorted in the ascending order of their maximum error lower bounds, obtained from logic simulation. The motivation is that only the top $K$ LACs are kept after the sorting. We want to keep as many valid LACs as possible within the top $K$ so that the circuit area can be effectively reduced by applying these valid LACs. Given that a LAC with a small maximum error lower bound is more likely to be valid, sorting the LACs by the lower bounds can increase the probability of keeping valid LACs in the top $K$ LACs.
Only if multiple LACs share the same maximum error lower bound, they are further sorted in descending order of the estimated area reduction, prioritizing LACs that reduce more area. Our method supports circuits represented in any graph format, such as AIGs and gate netlists. For AIGs, a LAC’s area reduction is estimated by counting the nodes in the maximum fanout-free cone (MFFC) [6] of the node where the LAC is applied, instead of performing time-consuming logic synthesis to obtain the exact reduction. For gate netlists, a LAC’s area reduction is obtained by summing the areas of the gates in the corresponding MFFC.
@@ V-A Experimental Setup
We implement the proposed ALS flow under the maximum error constraint in C++ and test it on a single core of an AMD Ryzen 9 5900X processor with 64GB RAM. Our flow is developed upon a state-of-the-art open-source logic synthesis and verification system, ABC [1]. The flow also integrates CryptoMiniSat [32], a modern SAT solver with rich features and a simple interface. To avoid long runtime of SAT solving, we set a maximum conflict number of $2^{18}$ for each SAT problem.
@@
In all experiments, the initial circuits are represented as AIGs, although our flow can also handle other circuit representations, such as gate netlists. The reason for using AIGs is that many AIG-based logic synthesis works [25, 3] and ALS works [19, 13] have shown significant advantages in reducing hardware cost, particularly for CMOS technologies. The standard cell library used in our experiments is the Nangate 45nm library [23]. The traditional logic synthesis in the last step of our flow (see Fig. 2) consists of AIG optimization and technology mapping. For AIG optimization, we apply the ABC script “resyn2rs” for 3 iterations. This script combines four transformations, i.e., rewriting, refactoring, balancing, and resubstitution, to simplify the local structure of the AIG, and has been shown effective and widely used in recent works [34, 7, 2, 36]. Iteratively applying the script can further reduce the AIG size and depth, and in our experiments, 3 iterations provide a good trade-off between circuit quality and runtime. For technology mapping, we use the standard area-oriented mapping script “dch; amap” to convert the AIG into a gate netlist, as recommended in the ABC tutorial [20]. Unless otherwise specified, the following default parameters are used in all experiments. For the simulation-based LAC pruning (see Section IV-B1), we first use $M_{\textit{small}}=2^{10}$ simulation patterns to quickly filter out LACs inducing large errors, and then use $M=2^{13}$ simulation patterns for a more fine-grained filtering. For the parameter $K$ in the sorting strategy in Section IV-C3, we set $K=100$. To build the error miter in Fig. 1 for maximum error checking, we use Verilog to describe the error miter and then use Yosys [43] and ABC to synthesize the Verilog description.
@@
To evaluate the hardware cost, we measure the area, delay, and power consumption of the synthesized gate netlist (post-synthesis, before place-and-route). The area is computed by summing the individual gate areas in the netlist. The delay is obtained using the static timing analysis command “stime” in ABC. The power is estimated with the Synopsys Design Compiler [35] at 2MHz under a uniform input distribution. We then define area ratio, delay ratio, and power ratio as the respective values of the approximate netlist normalized to those of the accurate one. Smaller ratios are preferred. To evaluate the accuracy of circuits, two different maximum error metrics, MaxED and MaxHD, are considered in our experiments. Note that the focus of this work is ALS under the maximum error constraint, so we do not compare it with other ALS methods under average error constraints, such as [13] and [15]. For all generated approximate circuits, the error miter in Fig. 1 is used to formally verify that the maximum errors of the circuits satisfy the given error bounds.
@@
TABLE I: Experimental benchmarks. Area, delay, and power are measured by mapping AIGs into gate netlists with the Nangate 45nm library.
@@
Size Depth Area $um^{2}$ Delay /$ns$ Power $\mu W$
Used in MECALS [17] absdiff 16/8 141 14 87.3 0.42 80.9
add8 16/9 66 10 42.0 0.36 36.8
add32 64/33 252 64 184.6 1.84 173.0
binsqrd 16/18 1562 50 1052.3 1.53 989.5
buttfly 32/34 265 48 170.5 1.01 172.6
mac 12/8 145 20 92.8 0.60 73.4
mult8 16/16 649 40 435.4 1.26 410.2
mult16 32/32 1981 72 1418.8 1.98 1707.2
mult32 64/64 8340 53 5723.3 1.87 7868.3
EPFL* arith- metic add128 256/129 1297 28 933.4 0.96 825.0
bar 135/128 2688 14 1267.8 0.92 1753.1
log2 32/32 38540 419 21480.6 14.26 40410.0
max 512/130 2686 549 1646.3 15.81 1971.2
mult64 128/128 33242 326 16447.3 9.47 31105.0
sine 24/25 7044 180 4112.1 5.93 5968.1
sqrt 128/64 21951 4591 13464.1 216.92 58397.0
square 64/128 20030 296 12801.8 7.96 18836.0
ISCAS85 c880 60/26 313 22 198.2 0.59 129.8
c1355 41/32 390 16 235.9 0.56 260.1
c1908 33/25 367 25 229.6 0.86 222.6
c2670 233/140 579 17 385.2 0.68 325.4
c3540 50/22 937 32 521.1 1.02 404.4
c5315 178/123 1306 28 720.3 0.72 643.7
c7552 207/108 1469 26 903.6 1.43 898.8